An object of mass 3 kg is moving on a rough surface with a velocity of 18m/s .It
covers a distance of a of 20m before coming to rest. find the the opposing force.
1
Expert's answer
2018-07-09T15:33:08-0400
Here, object mass m = 3 kg
Initial velocity (u) = 18 m/s
Final velocity (v) = 0 m/s
Distance travelled (S) = 20 m.
Now, we know, 〖 v〗^2=u^2-2fS ……………………..(1)
Where, f = deceleration of the object
Put the values of v , u and S in equation (1), we get,
0 = 〖18〗^2-2×f×20
Or, f = 〖18〗^2/40 = 8.1 m/s2.
Now , force = m × f = 3 × 8.1 = 24.3 N